OpenStudy (anonymous):

Find the half life of the following quantity; x(t)=90+10^(-4t)

OpenStudy (loser66):

what does half life mean?

OpenStudy (anonymous):

The time it takes for half of the substance to deteriorate.

OpenStudy (anonymous):

Would I set x(t)=180, then take log base 10 of each side?

OpenStudy (wolf1728):

I'm used to half life problems that are stated in English : 50 gms decays to 43 grams in 6.2 years - that sort of thing.

OpenStudy (anonymous):

They are usually stated in terms of exponential decay. To find the half life of y(t)=80*e^(-3t), for example, you would set y(t)=40. Then divide to get 1/2=e^(-3t). Then take the natural log of both sides to get -ln(2)=-3t, or t=ln(2)/3. I'm just not sure what to do when the equation is changed.

OpenStudy (tkhunny):

As stated, the problem makes no sense. 90 is constant. "half-life" is not a meaningful concept.

OpenStudy (wolf1728):

I'm usually very good at half life problems but I need to do some research on this. First of all in the equation x(t)=90+10^(-4t) what is the '90' and what does the '4' represent? I'm guessing the x(t) is the half life.

OpenStudy (anonymous):

That is all the information we are given. I'm assuming the -4 is the k, and 90 must be some sort of initial condition? I think x(t) is the half life as well.

OpenStudy (anonymous):

It doesn't seem to make too much sense as it is written.

OpenStudy (wolf1728):

I'll do some research. As I said I'm pretty good at half life problems. As a matter of fact, here is a page from my website: www.1728.org/halflife.htm

OpenStudy (wolf1728):

Hmmm you can't post a link in here?

OpenStudy (anonymous):

I just copied and pasted it into my browser, it worked.

OpenStudy (wolf1728):

Okay Melody :-) Time for me to do some research.

OpenStudy (anonymous):

Thank you! :)

OpenStudy (wolf1728):

Well I searched all over the internet and could not find any half life information that has a formula like this x(t)=90+10^(-4t) What a lousy way to phrase a question. I suppose you could tell this to your instructor. Well I tried. :-)

OpenStudy (anonymous):

Thank you for trying! I just e mailed them and asked if it was perhaps a mistake, or to see if they could offer any tips.

OpenStudy (anonymous):

I think it must have been a typo in the question.

OpenStudy (wolf1728):

Well you can always E-Mail them a question like " Since 60 miles per hour = 88 feet per second - when will the bus get to Philadelphia?"

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

I bet they would love that. Hopefully questions will be more clear in the future.

OpenStudy (wolf1728):

Yes, that might show them it is worth the time to state things as clearly as possible. Well if you do get an e-mail please let me know. see you Melody.

OpenStudy (anonymous):

I'll let you know their response!

OpenStudy (wolf1728):

thanks :-)

OpenStudy (anonymous):

The teacher just emailed me back, they changed the question to find the half life of 90*e^(-4*t). Much easier now!

OpenStudy (wolf1728):

I'm really interested in this now. I found a formula: P = A * e^(-time * ln(2)/half-life ) where p = present amount A = beginning amount and (ln(2)/half-life) is usually represented by 'k'. So I'm guessing that in the formula 90*e^(-4*t) '90' is the beginning amount 't' is the elapsed time and -4 is k (ln(2)/half-life) So, it seems to me that the question gives you the beginning amount and 'k' but to me wouldn't you also need the present amount to determine half-life ?